OpenStudy (anonymous):

(a3 - 2a + 5) - (4a3 - 5a2 + a - 2)

OpenStudy (jdoe0001):

\(\bf (a^3 - 2a + 5) - (4a^3 - 5a^2 + a - 2)\\ \quad \\ \textit{recall that } -\times -=+\quad and \quad +\times -=-\qquad thus\\ \quad \\ (a^3 - 2a + 5) - (4a^3 - 5a^2 + a - 2)\\ \quad \\ \implies a^3 - 2a + 5 - 4a^3 + 5a^2 - a + 2\\ \quad \\ \implies a^3 - 4a^3+ 5a^2- 2a - a + 2+ 5\)

OpenStudy (jdoe0001):

add up like-terms

OpenStudy (texaschic101):

(a^3 - 2a + 5) - (4a^3 - 5a^2 + a - 2) distribute the negative sign through the parenthesis a^3 - 2a + 5 - 4a^3 + 5a^2 - a + 2 put your like terms together... a^3 - 4a^3 = ? 5a^2 (there is no other like term with this one) -2a - a = ? 5 + 2 = ? once you get them together, then list them....what do you get ?